Directions: Solve each problem by using an absolute value equation or inequality.
In July 1985 Steve Cram of Great Britain set a world record of 3 minutes 29.67 seconds for the 1500 meter race and a world record of 3 minutes 46.31 seconds for the mile 1- mile race. Suppose you know only that these two events occurred 11 days apart and that the 1500 meter record was set on July 16. Use an absolute value equation to find the possible dates for the 1 mile record run.

Answer by stanbon(29486)

In July 1985 Steve Cram of Great Britain set a world record of 3 minutes 29.67 seconds for the 1500 meter race and a world record of 3 minutes 46.31 seconds for the mile 1- mile race. Suppose you know only that these two events occurred 11 days apart and that the 1500 meter record was set on July 16. Use an absolute value equation to find the possible dates for the 1 mile record run.
---------
The dates range from 11 days before July 16 to 11 days after July 16.
-----------------------
Draw a number line:
Plot the numbers 16-11=5, 16, and 16+11= 27
---------------------
Let x be any of the numbers from 5 to 27
----------------------------------
The distance from x to 16 is less than or equal to 11.
---------------------------------
In absolute language that means |x-16|<=11
=====================
Cheers,
Stan H.